one of the student asks the age of Ali his answer is: "when i was at your age you were 12 year old & when you are at my age ,i will be 66 year old"accordingly ,what is the student's current age?

Answer by ankor@dixie-net.com(22740)

one of the student asks the age of Ali his answer is:
"when i was at your age you were 12 year old & when you are at my age,i will be 66 year old "accordingly ,what is the student's current age?
:
I couldn't come up with a system of equations for this, Came up with this method
:
A line representing their present ages
12------------S------------A------------66
:
Find the difference: 66 - 12 = 54
Divide the difference by 3: 54/3 = 18 yrs is the difference in their ages
therefore:\
Student is 12 + 18 = 30 yrs
and
Ali is 30 + 18 = 48 yrs old
:
Check in the statement
"when i was at your age you were 12 year old,"
Ali: 48 - 18 = 30 which is the students present age
Student: 30 - 18 = 12 yrs
:
"when you are at my age,i will be 66 year old "
Student: 30 + 18 = 48 which is Ali's present age
Ali: 48 + 18 = 66 yrs